# Haldane model on the Sierpiński gasket

### v2

@@ V Effect of disorder on the topological states
In Fig. 12, we show the energy spectra of and some states on the Sierpiński gasket considered in Fig. 5 with this disorder for two different values of $\Delta$. We depict the six edge-localised, topological states from Fig. 5 (coloured) together with three bulk-like states (grey). In Fig. 12(a), the disorder parameter $\Delta=0.02\lambda$, and in (b) $\Delta=0.2\lambda$. For the smaller disorder, the edge states are virtually unaffected. For a disorder of about 20% of the NNN hopping, the topological states remain fully localised on the original edges, although the spatial symmetries are now clearly broken. This is similar to the behaviour observed for the quantum spin Hall effect exhibited by bismuth fractals on InSb [52]. This behaviour corroborates the topological nature of these edge-localised states (coloured). In contrast, the bulk states (grey), change their localisation significantly when the strength of the disorder is changed, because these states lack topological protection.
@@ VI Conclusions
To conclude, the Haldane model on a single Sierpiński gasket was considered. The energy spectra revealed the opening of flux-induced and fractal gaps. A staggered mass, which breaks the symmetry between the two sublattices, combined with a complex NNN hopping was shown to cause the formation of a set of topological states in the Sierpiński gasket. We investigated the robustness of these states against disorder and found that they remained fully edge-localised, indicative of their topological nature. To further characterise the topology of these states, we calculated the real-space Chern number, as proposed by Kitaev [58]. Using this method, we elucidated the relationship between the topological phases of the two types of gaps in these fractal structures and the different parameters of the model. Then, we compared the obtained phase diagram with previous results in the literature and showed that the Sierpiński gasket did not fit the original phase diagram presented by Haldane in two dimensions [4] or the squeezed phase diagram of a doubled Sierpiński gasket, as considered in Ref. [53]. By thoroughly investigating the topological phases arising in the flux-induced gaps and the largest fractal gap, we found that more intricate structures arise. These structures are reminiscent of Arnold’s tongues, suggestive of a possible connection to the fractality of this system. Our work lays a solid foundation for further studies of the enhancement of topological phases by fractality.
In the future, it would be interesting to extend this analysis to other topological models and/or other fractal structures. Furthermore, it could be interesting to investigate which role self-similarity plays and to make predictions for the thermodynamic limit of this fractal. Lastly, it would be interesting to experimentally realise the proposed topological phase transitions by tuning solely an applied perpendicular electric field.
@@ Appendix A Results of the void-respecting fractal Haldane model
Figure 13: The results of the considerations of the main text on a void-respecting Haldane model. (a)-(f) depict the Sierpiński cocoons of this void respecting Haldane model, analogous to Fig. 8. In (a)-(c), the energy $E$ corresponding to a gap is considered and in (d)-(f) the filling fraction $\eta$. In (g) and (h), energy spectra for a range of $\lambda$’s are presented. (g) corresponds to the SOC-like fractal Haldane model, and (h) to the void-respecting fractal Haldane model, as shown in the accompanying lattice of a second-generation fractal of each of these models. In (i), we show an energy spectrum, with the filling fractions of interest marked, and the corresponding phase diagrams are presented on a backdrop of matching colour in (j), akin to Fig. 11.
@@
In Figs. 13 (g) and (h), we present the energy spectra for different values of $\lambda$ (at $M=0$) for the SOC-like Haldane model of the main text (g), and the void-respecting Haldane model (h). These spectra are accompanied by a visual representation of the second generation of the fractal model considered, in which the NN and NNN hoppings are depicted in black (full) and grey (dashed) lines, respectively. Here, we observe how the inclusion of the complex NNN hoppings across the voids influences the dispersion of the spectra. Most notably, the emergence of a large flat band is observed for $\lambda=1/\sqrt{6}\approx 0.41$.
In Fig. 13(j), we present a set of phase diagrams of the system shown in Fig. 13(h), for different values of $\lambda$ and two different filling fractions $\eta=\{0.46,0.4\}$, analogous to Fig. 11. Here, we observe that the regions corresponding to a topological phase are both more stable and better quantized, but no increasingly complex patterns are observed. For the flux-induced gap (blue), we observe a similar trend as for the SOC-like Haldane model, where two (inverted) conical shapes form around $\Phi=0$ and $\Phi=\pm\pi$, reaching larger values of |$\absolutevalue{M}$ | with increasing $\lambda$. In addition, distinct upward (downward) phase boundaries (white curves) are shown in the centre of these conical shapes. For the fractal gap (purple), a similar conic shape arises, which reaches lower values of $M$ as $\lambda$ is increased, but no growing complexity is observed. Finally, in contrast with the SOC-like Haldane model, no droplet shapes emerge in the phase diagrams of the considered gaps of the void-respecting Haldane model.
@@ Appendix B Effect of the individual Haldane terms
Figure 14: The energy spectra of the SOC-like Haldane model in the bulk, a triangular flake, and the Sierpiński gasket. (a) The spectra for $\lambda=M=0t$, corresponding to Figs. 4(a)-(c) of the main text, (b) the spectra for $\lambda=0.3t$ and $M=0$, (c) for $\lambda=0t$ and $M=0.3t$, and (d) for $\lambda=M=0.3t$. This shows how either a nonzero Semenoff mass $M$ or a nonzero complex NNN hopping can open a bulk gap, but that the topological nature of these gaps differs. Furthermore, one may observe that the symmetry of the spectrum is lifted for the finite-size models for $M>0$, and how $\lambda>0$ induces the emergence of states crossing the gaps of the respective models.
@@
Appendix C Visualisation of the RSCN mask
@@
In the main text, we discussed how the value of the RSCN is not enough to conclusively decide about the nontriviality of a state. We emphasized that one also needs to establish whether such a state belongs to a “bulk band” or if such a state is “in-gap”. In this work, we took a numerical derivative and mapped the results to a value between 00 and $1$ to form a transparency mask by taking an exponential,
@@
In Fig. 15, we follow a similar procedure, but now the background of each figure is black, such that this mask actually mixes in a black colour. In this manner, not only the value but also the stability of the RSCN is depicted. Interestingly, while in most graphs just the boundaries of the topological phase are affected, in Fig. 15(b), four bright black and straight lines arise. These lines correspond to the bump in Fig. 6, which was discussed in the main text. This seems to suggest that there is some linear relationship between the strength of the staggered mass and the energy corresponding to this unstable divergence of the RSCN. This calls for further investigation, with alternative methods.
Appendix D Phase diagram for different generations
@@
In Fig. 16, the phase diagrams of both a flux-induced gap and the largest fractal gap of the Sierpiński gasket are shown. Here, the generation $g$ ranges from 3 to 6 and the complex NNN hopping strength $\lambda=0.8t$. In Fig. 16(a), the phase diagrams of the third-generation fractal are shown. Most notably, we see that the RSCN is poorly quantised and little patterns can be distinguished. This is because each subset $A$, $B$, and $C$ now contains about four sites, which is not large enough for the RSCN to confidently predict the topology of a gap.
Similarly, in Figs. 16(b), (c), and (d) the phase diagrams are shown for a fourth, fifth and sixth generation of the fractal, respectively. Here, we recognise two patterns: firstly, as the generation increases, the RSCN is better quantized to integer values, and secondly, with increasing generation the structure becomes more complex.
This first pattern is best observed for larger values of |$\absolutevalue{M}$ | in both phase diagrams, where the transition from |$\absolutevalue{\tilde{\nu}(P)}=1$ to zero and back becomes more stepwise, and the gradual transitions of the lower generations is slowly lost, caused by the better quantisation of the RSCN.

### v3

@@ V Effect of disorder on the topological states
In Fig. 12, we show the energy spectra of and some states on the Sierpiński gasket considered in Fig. 5 with this disorder for two different values of $\Delta$. We depict the six edge-localized, topological states from Fig. 5 (colored) together with three bulk-like states (gray). In Fig. 12(a), the disorder parameter $\Delta=0.02\lambda$, and in (b) $\Delta=0.2\lambda$. For the smaller disorder, the edge states are virtually unaffected. For a disorder of about 20% of the NNN hopping, the topological states remain fully localized on the original edges, although the spatial symmetries are now clearly broken. This is similar to the behavior observed for the quantum spin Hall effect exhibited by bismuth fractals on InSb [63]. This behavior corroborates the topological nature of these edge-localized states (colored). In contrast, the bulk states (gray), change their localization significantly when the strength of the disorder is changed, because these states lack topological protection.
@@ VI Conclusions
To conclude, the Haldane model on a single Sierpiński gasket was considered. The energy spectra revealed the opening of flux-induced and fractal gaps. A staggered mass, which breaks the symmetry between the two sublattices, combined with a complex NNN hopping was shown to cause the formation of a set of topological states in the Sierpiński gasket. We investigated the robustness of these states against disorder and found that they remained fully edge-localized, indicative of their topological nature. To further characterize the topology of these states, we calculated the real-space Chern number, as proposed by Kitaev [70]. Using this method, we elucidated the relationship between the topological phases of the two types of gaps in these fractal structures and the different parameters of the model. Then, we compared the obtained phase diagram with previous results in the literature and showed that the Sierpiński gasket did not fit the original phase diagram presented by Haldane in two dimensions [4] or the squeezed phase diagram of a doubled Sierpiński gasket, as considered in Ref. [64]. By thoroughly investigating the topological phases arising in the flux-induced gaps and the largest fractal gap, we found that more intricate structures arise. These structures are reminiscent of Arnold’s tongues, suggestive of a possible connection to the fractality of this system. Our work lays a solid foundation for further studies of the enhancement of topological phases by fractality.
In the future, it would be interesting to extend this analysis to other topological models and/or other fractal structures. Furthermore, it could be interesting to investigate which role self-similarity plays and to make predictions for the thermodynamic limit of this fractal. Lastly, it would be interesting to experimentally realize the proposed topological phase transitions by tuning solely an applied perpendicular electric field.
Figure 13: The results on a void-respecting Haldane model. (a)-(f) The Sierpiński cocoons of this void respecting Haldane model, in terms of energy $E$ [13(a)-(c)] and in terms of the filling fraction $\eta$ [13(d)-(f)]. In (g) and (h), energy spectra for a range of $\lambda$’s are presented. (g) corresponds to the SOC-like fractal Haldane model, and (h) to the void-respecting fractal Haldane model, as shown in the accompanying lattice of a second-generation fractal of each of these models. In (i), we show an energy spectrum, with the filling fractions of interest marked, and the corresponding phase diagrams are presented on a backdrop of matching color in (j), akin to Fig. 11.
@@ Appendix A Results of the void-respecting fractal Haldane model
In Figs. 13 (g) and (h), we present the energy spectra for different values of $\lambda$ (at $M=0$) for the SOC-like Haldane model of the main text (g), and the void-respecting Haldane model (h). These spectra are accompanied by a visual representation of the second generation of the fractal model considered, in which the NN and NNN hoppings are depicted in black (full) and gray (dashed) lines, respectively. Here, we observe how the inclusion of the complex NNN hoppings across the voids influences the dispersion of the spectra. Most notably, the emergence of a large flat band is observed for $\lambda=1/\sqrt{6}\approx 0.41$.
Figure 14: The energy spectra of the SOC-like Haldane model in the bulk, in a triangular flake, and in the Sierpiński gasket. (a) The spectra for $\lambda=M=0t$, corresponding to Figs. 4(a)-(c) of the main text, (b) the spectra for $\lambda=0.3t$ and $M=0$, (c) for $\lambda=0t$ and $M=0.3t$, and (d) for $\lambda=M=0.3t$. This shows that either a nonzero Semenoff mass $M$ or a nonzero complex NNN hopping can open a bulk gap, but that the topological nature of these gaps differs. Furthermore, one may observe that the symmetry of the spectrum is lifted for the finite-size models for $M>0$, and how $\lambda>0$ induces the emergence of states crossing the gaps of the respective models.
In Fig. 13(j), we present a set of phase diagrams of the system shown in Fig. 13(h), for different values of $\lambda$ and two different filling fractions $\eta=\{0.46,0.4\}$, analogous to Fig. 11. Here, we observe that the regions corresponding to a topological phase are both more stable and better quantized, but no increasingly complex patterns are observed. For the flux-induced gap (blue), we observe a similar trend as for the SOC-like Haldane model, where two (inverted) conical shapes form around $\Phi=0$ and $\Phi=\pm\pi$, reaching larger values of |$\absolutevalue{M}$ | with increasing $\lambda$. In addition, distinct upward (downward) phase boundaries (white curves) are shown in the center of these conical shapes. For the fractal gap (purple), a similar conic shape arises, which reaches lower values of $M$ as $\lambda$ is increased, but no growing complexity is observed. Finally, in contrast with the SOC-like Haldane model, no droplet shapes emerge in the phase diagrams of the considered gaps of the void-respecting Haldane model.
@@ Appendix B Effect of the individual Haldane terms
Figure 15: The energy spectra for different sizes of the triangular flake and the Sierpiński gasket, each analogous to Fig.4. The sizes are chosen to correspond to a generation $g=3,4,5$, respectively. The fractal gaps are marked in orange and the flux-induced gaps are marked in gray. These results allow us to distinguish between intrinsic and finite-size effects, such as the spurious gap opening in the spectrum of a triangular flake. Notably, the size and position of the fractal gaps do not depend on the generation, indicating that the gaps remain in the thermodynamic limit.
In Fig. 14(c), only a non-zero staggered mass is taken into account, $M=0.3t$. Now, the system is in a trivial phase, and a bulk band gap is observed. As expected, there are no in-gap states for the triangular flake. In addition, we observe that the middle flat band is shifted from $E=0$ to $E=M$. Similarly, the gaps of the Sierpiński gasket remain open, and the middle flat band is shifted to $E=M$. This raising of the energy of the central flat band of the OBC geometries breaks the spectral symmetry of these models. This is a consequence of $M\neq 0$, in combination with the asymmetry in the number of lattice sites in each sublattice. In Fig. 14(d), these effects are combined, and both the “in-gap” states and the broken symmetry can be observed.
Appendix C Investigation of finite-size effects
@@
Here, we investigate how the energy spectra discussed in Sec. III.1 scale with the size of the system. To this end, we have considered structures of sizes corresponding to generations $g=3,4,5$ in Fig. 15. For the smallest systems ($g=3$), a apparent gap around $E=0$ can be seen in the spectrum of a triangular flake. However, when we compare this spectrum to the spectrum of triangular flakes with more sites ($g=4$), the gap shrinks. For the largest considered system ($g=5$), this apparent gap has shrunk even further. Therefore, we conclude that this gap opening is a finite-size effect. This can be understood by a simple analogy with the particle-in-a-box model. In this model, the wavefunction must have nodes on the boundaries of the system. Consequently, the wavefunction with the lowest allowed frequency has frequency $f\approx(2L)^{-1}$, with $L$ a characteristic length scale. Therefore, the size of the triangular flake puts a lower limit on the allowed non-zero energies because the low-energy modes correspond to low-frequency wavefunctions.
It is of note that the fractal gaps behave very differently. The size and position of these gaps does not change when comparing different generations. Therefore, we conclude that these gaps are not a finite-size effect, but truly an intrinsic property caused by the fractality of the system.
Appendix D Visualization of the RSCN mask
In the main text, we discussed how the value of the RSCN is not enough to conclusively decide about the non-triviality of a state. We emphasized that one also needs to establish whether such a state belongs to a “bulk band” or if such a state is “in-gap”. In this work, we took a numerical derivative and mapped the results to a value between 00 and $1$ to form a transparency mask by taking an exponential,
@@ Appendix D Visualization of the RSCN mask
In Fig. 16, we follow a similar procedure, but now the background of each figure is black, such that this mask actually mixes in a black color. In this manner, not only the value but also the stability of the RSCN is depicted. Interestingly, while in most graphs just the boundaries of the topological phase are affected, in Fig. 16(b), four bright black and straight lines arise. These lines correspond to the bump in Fig. 6, which was discussed in the main text. This seems to suggest that there is some linear relationship between the strength of the staggered mass and the energy corresponding to this unstable divergence of the RSCN. This calls for further investigation, with alternative methods.
Appendix E Phase diagram for different generations
@@
In Fig. 17, the phase diagrams of both a flux-induced gap and the largest fractal gap of the Sierpiński gasket are shown. Here, the generation $g$ ranges from 3 to 6 and the complex NNN hopping strength $\lambda=0.8t$. In Fig. 17(a), the phase diagrams of the third-generation fractal are shown. Most notably, we see that the RSCN is poorly quantized and little patterns can be distinguished. This is because each subset $A$, $B$, and $C$ now contains about four sites, which is not large enough for the RSCN to confidently predict the topology of a gap.
Similarly, in Figs. 17(b), (c), and (d) the phase diagrams are shown for a fourth, fifth and sixth generation of the fractal, respectively. Here, we recognize two patterns: firstly, as the generation increases, the RSCN is better quantized to integer values, and secondly, with increasing generation the structure becomes more complex.
This first pattern is best observed for larger values of |$\absolutevalue{M}$ | in both phase diagrams, where the transition from |$\absolutevalue{\tilde{\nu}(P)}=1$ to zero and back becomes more stepwise, and the gradual transitions of the lower generations is slowly lost, caused by the better quantization of the RSCN.